A 10 inch T.v. set has a width of 8 inches, so what is the height?
Answer by npinson90(1)

I would use the pythagorean theorem to solve this problem.
Since televisions are measured diagonally, you have a right triangle.
Our hypotenuse is 10 and we will consider the width {b=8}.
Our unknown is a.
(a^2)+(b^2)=(c^2)
So we have (a^2)+(8^2)=(10^2)
(a^2)+64=100
(a^2)+64-64=100-64
(a^2)=36
Now we will take the square root of 36, and it happens to
be a perfect square so there are 2 solutions = (-6, 6)
Since -6 is not a possible length, we will throw out that solution.
Answer, the height is 6 inches.